OpenStudy (anonymous):

6 sqrt 72x^2 show work

OpenStudy (mathstudent55):

Is the question \(\large 6\sqrt{72x^2}\) ?

OpenStudy (mathstudent55):

Are you told anything about x, such as assume all variables are non-zero?

OpenStudy (mathstudent55):

@MeganChase Hint: 72 = 36 * 2 In other words, you can take out more than just sqrt(9) from the radical.

OpenStudy (anonymous):

i forgot to go higher. i suck at math. i need to listen more in class. bye. time to hit the books.

OpenStudy (mathstudent55):

Then there is the question iof what to do with \(\sqrt{x^2} \). If no restrictions on x are given, then \(\sqrt{x^2} = |x|\).

OpenStudy (mathstudent55):

^of

OpenStudy (anonymous):

im trying to simplify the radical

OpenStudy (mathstudent55):

\(\large 6\sqrt{72x^2}\) \(\large = 6\times \sqrt{36 \times 2 \times x^2}\) \(\large = 6\times \sqrt{36} \times \sqrt{2} \times \sqrt{x^2}\) \(\large = 6\times \color{red}{\sqrt{36}} \times \color{green}{\sqrt{x^2}} \times \color{blue}{\sqrt{2}}\) \(\large = 6 \times \color{red}{6} \times \color{green}{|{x|}} \times \color{blue}{\sqrt{2}}\) \(\large =36\sqrt{2}|x| \)

OpenStudy (mathstudent55):

GTG. If you have any questions, post them. I'll try to answer when I get back online.

OpenStudy (anonymous):

okay and thank u